40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 29 19:38:09 2015, gautam, Update, General, FOL fiber box revamp IMG_0017.JPGFOL_schematic.pdfIMG_0018.JPGIMG_0016.JPG
    Reply  Wed Sep 30 13:07:13 2015, gautam, Update, General, FOL fiber box revamp FOL_schematic.pdf
       Reply  Tue Oct 6 16:56:40 2015, gautam, Update, General, FOL fiber box revamp 20151006_Xbeat_psd.pdf20151006_Xbeat_coherence.pdf
Message ID: 11652     Entry time: Wed Sep 30 13:07:13 2015     In reply to: 11650     Reply to this: 11670
Author: gautam 
Type: Update 
Category: General 
Subject: FOL fiber box revamp 

Eric pointed out that the 1x2 couplers that were used in the previous arrangement and which I recycled, were in fact NOT appropriate - they are not 50-50 couplers but 90-10 couplers, which explains the measured power levels I quoted here.

I switched out these for a pair of the newly arrived 2x2 couplers, and have also replaced the datasheets on the inside of the top cover. I then redid the power level measurements, and got some sensible values this time (see Attachment #1 for revised layout and measured power levels, numbers in red are powers for PSL light, numbers in green are for AUX laser light, and all numbers are in mW). I did find that the 90-10 splitter in the PSL+Y path was not working (though the one in the PSL+X path seems to be working fine), and hence, have not quoted power levels at the output of these splitters. For now, I guess we can bypass the splitters and take the PSL+AUX light from the 2x2 couplers directly to the PDs.

Attachment 1: FOL_schematic.pdf  61 kB  | Hide | Hide all
FOL_schematic.pdf
ELOG V3.1.3-